### v1

# How many qubits does a machine learning problem require?

Abstract
For a machine learning paradigm to be generally applicable, it should have the property of universal approximation, that is, it should be able to approximate any target function to any desired degree of accuracy. In variational quantum machine learning, the class of functions that can be learned depend on both the data encoding scheme as well as the architecture of the optimizable part of the model. Here, we show that the property of universal approximation is constructively and efficiently realized by the recently proposed bit-bit encoding scheme. Further, we show that this construction allows us to calculate the number of qubits required to solve a learning problem on a dataset to a target accuracy, giving rise to the first resource estimation framework for variational quantum machine learning. We apply bit-bit encoding to a number of medium-sized datasets from OpenML and find that they require only $20$ qubits on average for encoding. Further, we extend the basic bit-bit encoding scheme to one that can handle batching very large datasets. As a demonstration, we apply this new scheme to the giga-scale transcriptomic Tahoe-100M dataset, concluding that the number of qubits required for encoding it lies beyond classical simulation capabilities. Remarkably, we find that the number of qubits does not necessarily increase with the number of features of a dataset, but may sometimes even decrease.
1 Introduction
The use of quantum computers for machine learning and artificial intelligence applications is a relatively new research area that could provide a means to model datasets which contain correlations that are hard to model classically. Quantum machine learning algorithms have been tested on near-term quantum hardware for various real-world datasets [1, 2, 3, 4, 5, 6, 7, 8, 9]. In particular, variational quantum machine learning, which uses parameterized unitaries that can be trained to model the correlations in a dataset analogous to classical neural networks [10], can be applied to a wide variety of problems. However, compared to other applications such as chemistry [11, 12, 13, 14] or cryptography [15], where algorithms with well-understood quantum advantage and extensively studied resource estimation are known, there are many more open questions around how to use quantum computers effectively for learning from data.
Among the fundamental questions is one of expressivity - how does one design a variational quantum ansatz, which we refer to as a quantum model, that can approximate the underlying function defined by a given dataset. To answer this question meaningfully, we must note that real-world datasets seldom follow theoretically tractable rules. Therefore, the quantum model should have the property of universal approximation - that is, it should be able to approximate any function with any desired degree of accuracy. Further, the construction which realizes the universal approximation should not itself require exponentially scaling computation, either quantum or classical. If a quantum learning model does not have this property, it is likely not useful in real life. Even if such a model works in a limited setting, we cannot expect it to give good results in the limit of large and complex datasets. Indeed, in classical machine learning, universal approximation theorems [16, 17] provide the mathematical justification for using neural networks for large and complex datasets.
[18] showed that the expressivity of a quantum model is constrained by the data-encoding method used. Specifically, a quantum model can be expanded as a partial Fourier series in the encoded data where the frequencies present in the series are set by linear combinations of the eigenvalues of the encoding operators. Following this, it is possible to increase the number of frequencies in the series by uploading the data multiple times, a technique known as data reuploading. The time for loading is then lower-bounded by the quantity of information to be loaded multiplied by the number of times it is loaded. With this method, even an image from a simple dataset like MNIST, which has 784 pixels, is expensive to load directly into a quantum state. Since this limit is fundamental, it is clear that some kind of compression is required before loading, for which techniques like principal component analysis are often used. However, it remains unclear how much the degree of compression and the number of data reuploadings affect the accuracy that can be reached by the model.
Here we show that the bit-bit encoding technique [19], which was recently proposed for supervised learning, directly addresses the problem of achieving efficient compression with predictable accuracy. In this encoding, both the input and the output are encoded as binary strings following a dimensionality reduction procedure. These binary strings correspond to the basis states of a quantum computer at initialization and measurement. For a small number of qubits, there are collisions in which the same input corresponds to more than one output, but as the number of qubits increases, the encoding becomes more precise. We use the Solovay-Kitaev theorem [20] to prove that a variational quantum model with bit-bit encoding has the property of universal approximation. Further, the classical pre-processing required before data loading utilizes techniques such as principal component analysis which scale only polynomially in the dataset size.
A very useful emergent property of the bit-bit encoding construction is that it is possible to efficiently calculate exactly the number of qubits required to encode a dataset to a desired degree of accuracy. This, then gives us the first framework for resource estimation in quantum machine learning. Using this framework, we can for the first time estimate the number of qubits required to solve a learning problem using a variational approach. In turn, this analysis, has implications for assessing the potential of quantum advantage for a dataset as we can calculate whether or not a learning problem fits into a number of qubits that can be simulated on a classical computer.
We introduce $Q_{\text{dataset}}$, a resource metric that estimates the number of qubits required to model a dataset using bit-bit encoding. Starting with the MNIST dataset, we illustrate how $Q_{\text{dataset}}$ is derived by incrementally increasing qubit allocations until collisions on both the train and test datasets vanish, revealing a trade-off between maximizing theoretical training accuracy and achieving generalization. Extending this analysis to a variety of OpenML datasets, we find that typical medium-sized classical machine learning datasets require $20$ qubits for coverage on average, suggesting they are unlikely to be candidates for quantum advantage [21]. Comparisons across dimensionality reduction methods show that schemes producing independent or de-correlated features achieve lower $Q_{\text{dataset}}$ than non-linear or unprocessed data. Finally, scaling experiments on the large Tahoe-100M transcriptomic dataset [22] reveal examples where quantum models require a number of qubits well beyond classical simulation capabilities, which peak at about 50 qubits [21], indicating that this is a good arena to look for quantum advantage.
We note that we only consider classical datasets in this work. Although there has been progress in quantum machine learning for quantum data [23, 24], its practical applicability remains limited. In contrast, quantum machine learning applied to classical data offers a far broader range of potential applications..
The paper is organized as follows. We begin with an overview of bit-bit encoding in Section 2, followed by an analysis of expressive power and universal approximation in commonly used encoding schemes in Section 3. Section 4 discusses several techniques that extend the original data encoding techniques proposed in paper [19], including techniques that can be used with large datasets. Section 4 also presents the number of qubits required for common datasets from OpenML. Section 5 demonstrates training a classifier on a bit-bit encoded dataset. We conclude with a discussion in Section 6.
2 Overview of Bit-Bit Encoding
Bit-bit encoding [19] encodes datasets with numeric features and discrete class labels into bitstrings, which allows the dataset to be read into a Hillbert space of a given number of qubits for classification. Consider a supervised learning problem in which we are given a dataset with consists of a feature matrix $X\in\mathbb{R}^{s\times n}$ and class label vector $y\in\{0,1,\ldots,c\}^{s}$, where $s$ is the number of samples, $n$ is the number of features, and $c$ is the number of class labels. For a classification problem, the learning task consists of finding an efficiently computable function that maps $X$ to $y$. Typically, the function is many-to-one, that is, $c<|X|$.
Bit-bit encoding converts the feature matrix $X$ into a discretized feature matrix $Z$, compactly encoded into $N_{x}$ qubits. The binary encoding algorithm works as follows. First, the dataset is split into train and test sets: $X_{\text{train}}$, $y_{\text{train}}$, $X_{\text{test}}$, and $y_{\text{test}}$. A dimensionality reduction model, principal component analysis (PCA) by default, is fit on $X_{\text{train}}$ and applied to $X_{\text{test}}$, reducing both sets to $D$ components. The number of bits assigned to each component is calculated with
$b_{d}=\lfloor N_{x}\cdot\frac{I(X_{d};y)}{\sum_{i=1}^{D}{I(X_{i};y)}}\rceil$
where $I$ is the mutual information and $N_{x}$ is the number of bits allocated to the discretized feature matrix. The dimension-reduced $X_{\text{train}}$ and $X_{\text{test}}$ are then min-max normalized so that each sample lies in $0,1]^{D}$. A copula transformation is fit on the reduced and normalized $X_{\text{train}}$ and applied to the reduced and normalized $X_{\text{test}}$, mapping each component to its empirical CDF so that its marginal distribution is approximately uniform within each class. Finally, to encode the reduced, normalized, and transformed $X_{\text{train}}$ and $X_{\text{test}}$, each data point $x_{sd}$ (the $s^{th}$ sample of the $d^{th}$ feature) is discretized into $b_{d}$ bits of precision with
$Z_{sd}=\lfloor x_{sd}2^{b_{d}}\rfloor$
where $Z$ is the discretized feature matrix. As the number of qubits $N_{x}$ grows, $Z$ provides a more precise representation of the original dataset.
Discretizing the feature matrix also compresses it, potentially creating collisions at the input as a particular feature vector $z\in Z$ may occur more than once, each time corresponding to the same or different value of $y_{s}\in y$. The samples in the discretized feature matrix can then be considered as samples from a joint probability distribution $f(Z=z,y=y_{s})$ over correlated random variables $(Z,y)$. For the purpose of classification, the ‘correct’ output is then taken to be the mapping which occurs more frequently, that is
$C(z)=\arg\max_{y_{s}}f(z,y_{s})$
[…]
The quantum learning problem can now be cast in terms of an operator $U_{*}$ that has the following action on $N_{q}=N_{x}+N_{y}$ qubits:
[…]
find the value of $\vec{\theta}$ which minimizes the loss function
$\bar{L}(\vec{\theta})=1-\sum_{z\in\mathcal{D}_{\vec{b}}}f(z)P_{C(z),z}(\vec{\theta}),$ (3)
where $f(z)$ is the frequency of occurrence of $z$ in the dataset. $0\leq\bar{L}\leq 1$, and has the simple interpretation of being the probability of the model giving the wrong answer when queried. If $U(\vec{\theta})=U_{*}$ and there is no noise, even one shot would suffice to classify the input data sample.
Note that the loss function only includes the action of $U$ on values of $z$ in the dataset. The action of $U$ on values of $z$ not in the dataset will determine its generalization behavior. In that case, a quantum model that is perfectly trained on the dataset will output a probability distribution over the possible classes.
We distinguish the bit-bit encoding technique from a simple basis encoding as it can be used to represent datasets that do not originally have a binary representation, but are converted into binary representation through a process of dimensional reduction followed by mutual-information based discretization. Indeed, a basis encoding that utilized only a naive discretization of the dataset would require many more qubits. For instance, a dataset consisting of molecular fingerprint samples [25], which are naturally binary, would take thousands of qubits to encode.
3 Universal Approximation
Next, we examine the expressive ability of bit-bit encoded quantum models and make a comparison with other encoding schemes. In this section, we discuss universal approximation for quantum machine learning and show that it can be straightforwardly achieved with bit-bit encoding whereas it is harder to construct efficiently for amplitude and angle encoding, the two other commonly used encoding schemes.
Definition: Let $\mathcal{X}\subseteq\mathbb{R}^{n}$ and $\mathcal{Y}\subseteq\mathbb{R}^{k}$ be compact sets. Consider a family of parameterized quantum models
$\mathcal{F}=\left\{f_{\theta}:\mathcal{X}\to\mathcal{Y}\,\middle|\,f_{\theta}(x)=\operatorname{Tr}\!\left(M\,U(\theta,x)\,\rho_{\mathrm{in}}\,U(\theta,x^{\dagger}\right),\;\theta\in\Theta\subseteq\mathbb{R}^{p}\right\},$
generated by preparing an input state $\rho_{\mathrm{in}}$, applying a parameterized quantum circuit $U(\theta,x)$, and measuring with a POVM $M$.
We say that $\mathcal{F}$ is universally approximating if, for every continuous function $g:\mathcal{X}\to\mathcal{Y}$ and every $\varepsilon>0$, there exists $\theta\in\Theta$ such that
$\sup_{x\in\mathcal{X}}\,\left\|f_{\theta}(x)-g(x)\right\|<\varepsilon.$
3.1 Universality of bit-bit encoding
In this section, we show how the bit-bit encoding leads to a universal approximation theorem for classification. First, we prove that the classification operator $U_{*}$ defined in Eq. 1 is unitary.
Theorem 1: Let $N_{x},N_{y}\in\mathbb{N}$, and set
$\mathcal{H}_{x}:=(\mathbb{C}^{2})^{\otimes N_{x}},\qquad\mathcal{H}_{y}:=(\mathbb{C}^{2})^{\otimes N_{y}},\qquad\mathcal{H}:=\mathcal{H}_{y}\otimes\mathcal{H}_{x}.$
[…]
$C(z)=C(z^{\prime})\;\Longrightarrow\;\braket{g(z)|g(z^{\prime})}=\delta_{z,z^{\prime}}.$ (4)
Define a linear map $V$ on the subspace
[…]
by
[…]
Then there exists a unitary $U_{*}:\mathcal{H}\to\mathcal{H}$ such that
[…]
Proof: First, for any $z,z^{\prime}\in\{0,1\}^{N_{x}}$,
$\bigl{\langle}C(z),g(z)\bigm{|}C(z^{\prime}),g(z^{\prime})\bigr{\rangle}$ $=\braket{C(z)|C(z^{\prime})}\,\braket{g(z)|g(z^{\prime})}$
$=\delta_{C(z),C(z^{\prime})}\,\braket{g(z)|g(z^{\prime})}.$
If $C(z)\neq C(z^{\prime})$ the first factor vanishes; if $C(z)=C(z^{\prime})$, the condition (4) gives $\braket{g(z)|g(z^{\prime})}=\delta_{z,z^{\prime}}$. Hence, in all cases,
$\bigl{\langle}C(z),g(z)\bigm{|}C(z^{\prime}),g(z^{\prime})\bigr{\rangle}=\delta_{z,z^{\prime}}.$ (7)
Therefore the set
[…]
is an orthonormal family in $\mathcal{H}$. Because $|\{0,1\}^{N_{x}}|=2^{N_{x}}$ and $\dim\mathcal{H}_{\mathrm{in}}=2^{N_{x}}$, we have $\dim\mathrm{span}(\mathcal{B}_{\mathrm{out}})=2^{N_{x}}$. Thus $V$ in (5) is an isometry from $\mathcal{H}_{\mathrm{in}}$ onto
$\mathcal{H}_{\mathrm{out}}:=\mathrm{span}(\mathcal{B}_{\mathrm{out}}),$
and, because these subspaces have equal dimension, $V$ is unitary between them.
[…]
Next, we show that the classification operator $U_{*}$ can be approximated to any desired degree of accuracy by a quantum circuit constructed out of a finite universal gate set.
[…]
Let $\mathcal{G}$ be any finite, fixed, two-qubit universal gate set. Then, for every $\varepsilon>0$, there exists a $\mathcal{G}$-circuit $U_{\varepsilon}$ on $N_{q}$ qubits such that
$\|U_{\varepsilon}-U_{*}\|_{\mathrm{op}}\;\leq\;\varepsilon,$ (9)
and, consequently, for every $z\in\{0,1\}^{N_{x}}$,
[…]
In particular, if the first $N_{y}$ qubits are measured in the computational basis, the outcome distribution under $U_{\varepsilon}$ differs from the ideal one by total variation distance at most $O(\varepsilon)$; hence the implemented classifier coincides with $C$ on all inputs $z$ except with probability $O(\varepsilon)$.
Proof: The existence of $U_{*}$ with the action (8) follows from Theorem 1.
[…]
The two theorems together show that the bit-bit encoding combined with an ansatz that consists of layers of gates from a universal gate set can approximate a classifier to any desired degree of accuracy on a dataset discretized as shown in the previous section. Finally, we observe that there exists a maximum number of bits required to encode the samples in a dataset.
Observation: Let $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{s}$ denote a dataset, where each feature vector $x_{i}$ admits a representation using at most $p_{x}$ bits, and each label $y_{i}$ admits a representation using at most $p_{y}$ bits. Then each pair $(x_{i},y_{i})$ can be uniquely encoded using no more than $p_{x}+p_{y}$ bits.
In fact, the limit where the above observation holds corresponds to a simple basis encoding of the dataset. Further, it is easy to see that as the number of bits increases, the accuracy of the encoding monotonically increases. Thus, as the number of qubits and depth of the ansatz increase, the quantum model can approximate the underlying function expressed by the dataset to any desired degree of accuracy, achieving universal approximation.
3.2 Other encoding schemes
Two other encoding schemes are often utilized in the quantum machine learning literature - amplitude encoding, and angle encoding. For both schemes, some kind of compression is required in order to make the data loadable into the quantum computer. Further, within each scheme, the data can be loaded into the quantum computer multiple times in order to enhance the model’s expressivity. However, the combined effects of the compression, data reuploading and design of the variational ansatz on the accuracy that can be attained by the model are not yet understood.
Amplitude encoding: A real input vector $x\in\mathbb{R}^{n}$ with $\|x\|_{2}=1$ under amplitude encoding is mapped to the quantum state $|x\rangle=\sum_{i=0}^{n-1}x_{i}\,|i\rangle\in\mathbb{C}^{n}$. Applying a parameterized unitary $U(\theta)$ yields the state $|\psi(x,\theta)\rangle=U(\theta)|x\rangle$. Measuring an observable $M$ produces the model output $f(x;\theta)=\langle\psi(x,\theta)|M|\psi(x,\theta)\rangle=\langle x|U^{\dagger}(\theta)MU(\theta)|x\rangle$. Defining $A(\theta):=U^{\dagger}(\theta)MU(\theta)$, we obtain $f(x;\theta)=x^{\top}A(\theta)x$ with $A(\theta)$ Hermitian. Hence the class of functions implementable by amplitude encoding followed by a parameterized unitary and measurement is
$\mathcal{F}_{\mathrm{AE}}=\bigl{\{}f(x)=x^{\top}A\,x:A\in\mathrm{Herm}(n),\ \|x\|_{2}=1\bigr{\}},$
namely the set of real quadratic forms on the unit sphere in $\mathbb{R}^{n}$. However, $C(S^{n-1})$, the space of continuous functions on the input sphere, is infinite-dimensional, whereas $\mathcal{F}_{\mathrm{AE}}$ is finite-dimensional. Hence $\mathcal{F}_{\mathrm{AE}}$ cannot be dense in $C(S^{n-1})$, meaning amplitude encoding does not satisfy universal approximation.
The expressivity of amplitude encoding can be enhanced by loading multiple states in parallel. However, a significant barrier remains in the form of the loading time being proporational to the dimension of the data. Previous work has shown the effectiveness of using tensor network techniques to compress the data before loading it as an amplitude encoded state [26], [4], [27], [28], but has not studied the effect of the compression on the model accuracy.
Angle encoding: Under angle embedding, each input $x\in\mathbb{R}^{n}$ is mapped to a product of single-qubit rotations $U_{\mathrm{enc}}(x)=\bigotimes_{j=1}^{d}R_{\alpha}(\phi_{j}(x))$ (where $R_{\alpha}$ are Pauli rotations), after which a parametrized ansatz $W(\theta)$ is applied and an observable $M$ is measured, yielding $f(x;\theta)=\langle 0|\,U_{\mathrm{enc}}^{\dagger}(x)W^{\dagger}(\theta)MW(\theta)U_{\mathrm{enc}}(x)\,|0\rangle$. Writing $A(\theta)=W^{\dagger}(\theta)MW(\theta)$ and expanding conjugation by $U_{\mathrm{enc}}(x)$ shows that $f$ is a finite trigonometric Fourier series in $x$, namely $f(x;\theta)=\sum_{k}a_{k}(\theta)\cos(\omega_{k}\cdot x)+b_{k}(\theta)\sin(\omega_{k}\cdot x)$ for some frequencies $\omega_{k}\in\mathbb{Z}^{n}$ arising from the Pauli rotation exponents. With $L$ layers of data reuploading, $|\psi(x,\theta)\rangle=\bigl{(}\prod_{\ell=1}^{L}W_{\ell}(\theta_{\ell})U_{\mathrm{enc}}(x)\bigr{)}|0\rangle$, the attainable frequency set expands to $\Omega_{L}\subset\mathbb{Z}^{n}$, so that the expressible function class is the family of trigonometric polynomials $\mathcal{F}_{\mathrm{angle}}^{(L)}=\bigl{\{}\,f(x)=\sum_{\omega\in\Omega_{L}}[a_{\omega}\cos(\omega\cdot x)+b_{\omega}\sin(\omega\cdot x)]:a_{\omega},b_{\omega}\in\mathbb{R}\,\bigr{\}}$.
The set of achievable frequencies $\Omega_{L}$ generated by nested Pauli rotations depends intricately on the circuit architecture and grows in a highly non-regular, combinatorial fashion. Consequently, there is no efficient algorithm that, given a dataset $\mathcal{D}$ or a target function $f^{\star}$, determines the minimum number of reuploading layers $L$ for which $\mathcal{F}_{\mathrm{angle}}^{(L)}$ contains an $\varepsilon$-approximant to $f^{\star}$. Moreover, unlike classical Fourier approximation, where one systematically improves approximation quality by increasing the maximum frequency or degree in a controlled way, the sets $\Omega_{L}$ are not nested in a structured fashion (for example, $\Omega_{L+1}$ does not simply add all frequencies above a certain cutoff), so increasing $L$ does not provide a principled or monotonic method of reducing approximation error. Instead, $L$ must be heuristically chosen, trained, and empirically validated, meaning the universal approximation property does not translate into a systematic, constructive approximation procedure.
In line with this, there are several recent papers that aim to enhance the expressivity of angle encoding. [29] adds auxiliary qubits to enhance the frequency spectrum of an architecture with angle embedding. [30] finds that using data reuploading with circuits that are wider rather than deeper is a more effective strategy. Despite this work on enhancing expressivity, there remains no known method for efficiently and constructively realizing universal approximation with these encoding schemes.
Combining these results, we conclude that bit-bit encoding provides a direct and efficient route to universality in quantum machine learning models. By construction, every dataset of finite precision can be uniquely represented using a bounded number of qubits, and the corresponding classifier can be realized as a unitary operator acting on the joint input–output space. The Solovay–Kitaev theorem then ensures that such unitaries can be approximated to arbitrary accuracy using a finite universal gate set, thereby establishing universality for bit-bit encoded models. In contrast, amplitude and angle encoding require additional structural assumptions or circuit resources in order to achieve comparable universality. Thus, bit-bit encoding yields both a conceptually transparent and practically scalable framework for constructing universally approximating quantum models.
4 Techniques: Bit Encoding Schemes
Bit-bit encoding maps numeric features and class labels into bitstrings for quantum classification. A necessary aspect of this process is compressing the original dataset. In this section, we empirically investigate how many qubits are required to fully cover (represent) bit-bit encoded machine learning datasets. We consider training and testing sets separately when calculating coverage. We calculate the training set coverage as 1training collision incidence1-$\text${training collision incidence}, where the training collision incidence is the proportion of training samples with different class labels which collide (i.e. have the same encoding). A training set is fully covered when there are no collisions between samples with different class labels (i.e. samples with the same encoding have the same class label). In this case, since there are no colliding samples with different classes, a perfect model would theoretically be able to achieve $100\%$ training accuracy. Therefore, we refer to 1training collision incidence1-$\text${training collision incidence} as the theoretical training accuracy. We define the testing set coverage as 1testing overlap incidence1-$\text${testing overlap incidence}, where the testing overlap incidence is the proportion of testing samples which overlap with the training data but do not have the same class label as the majority class label of the overlapping training data. A testing set is fully covered when all of the samples either 1.) are not overlapping with the training data or 2.) have the same class label as the majority class label of the overlapping training data. We refer to 1testing overlap incidence1-$\text${testing overlap incidence} as the theoretical testing accuracy. The minimum number of qubits at which the training and testing sets are covered (i.e. $100\%$ theoretical accuracy) is $Q_{\text{train}}$ and $Q_{\text{test}}$, respectively. The number of qubits required to cover the class labels is $Q_{y}=\lceil\log_{2}{(\text{number of classes})}\rceil$. The minimum number of qubits to cover an entire dataset is $Q_{\text{dataset}}=\max{(Q_{\text{train}},Q_{\text{test}})}+Q_{y}$. While in the bit-bit encoding algorithm $D$, the number of components after the dimensionality reduction scheme acts on the original features, is a modifiable parameter, in all experiments we set $D$ equal to the original number of features.
4.1 Simple PCA
Figure 1: The train and test theoretical accuracy (left), test-train overlap (middle), and correctly classified test-train overlap (right) over the number of allocated qubits for the MNIST $784$ OpenML dataset with PCA dimensionality reduction. Each line shows the mean of the corresponding metric with $95\%$ confidence interval error bands. The gray dotted line is the mean $Q_{\text{dataset}}$ (minus the class qubits) when the theoretical accuracy threshold is $0.99$, instead of the default $1.0$.
We start with a simple example of calculating $Q_{\text{dataset}}$ for the MNIST digit-recognition dataset. We use the MNIST $784$ dataset from OpenML [31], which converts the 2828x2828 images into $784$ features, where the features are the grayscale intensity of each pixel. The MNIST dataset contains $70,000$ samples and $10$ classes, one class for each single-digit number. Figure 1, left, shows the mean theoretical accuracy across $10$ replicates of training/testing splits. $Q_{\text{dataset}}$ is the number of class qubits plus the number at which both the training and testing theoretical accuracy reach $1$. The maximum $Q_{\text{dataset}}$ for MNIST $784$ is $63$, where the x-axis of the figure is cut-off, while the minimum $Q_{\text{dataset}}$ is $42$ and the mean is $50$. The gray dotted line shows the mean $Q_{\text{dataset}}$ ($27$) minus the number of class qubits ($4$) if the theoretical accuracy threshold was $0.99$ instead of $1.0$. The difference between $Q_{\text{dataset}}$ between thresholds is explained by the theoretical training accuracy slowly increasing from $0.99$ to $1.0$ over the number of qubits. We postulate that the difference between the $Q_{\text{dataset}}$ at the threshold of $0.99$ versus $1.0$ reveals a trade-off between achieving the maximum theoretical training accuracy and obtaining a generalizable testing set; as the number of qubits increases, the overlap between the training and test sets decreases (Figure 1, middle). Furthermore, as seen in Figure 1, right, the dip in the testing accuracy between $10$ and $20$ qubits is caused by shifts in how the unique bit strings overlap across classes as the encoding changes with the number of qubits allocated. Between $20$ and $50$ qubits, test-train overlap is near but not quite 0, which is why there is stochasticity in correctly classified test-train overlap in that range.
Next, we calculate $Q_{\text{dataset}}$ for a variety of benchmark classification datasets from OpenML [31]. We use all OpenML datasets that fulfill the following criteria: less than $1,000$ features which are all numeric, less than $10,000$ samples, more samples than features, no missing values, single-label classification with less than $10$ classes, and no collisions when not encoded. After filtering, we have $29$ datasets. Supplementary Figure S1 visualizes the number of samples, number of features, and number of class labels of each dataset. We randomly split each dataset into $80\%$ training and $20\%$ testing, repeated across $10$ replicates. We obtain $Q_{\text{dataset}}$ of each dataset by calculating the training collision and testing overlap incidence when continuously increasing the number of qubits allocated to the dataset from $1$ until the training and testing collision are both 0. We find the mean $Q_{\text{dataset}}$ for the OpenML datasets to be $20.49$ qubits, $95\%$ CI $[19.97,21.00]$. Supplementary Figure 2 shows the theoretical accuracy for the training and testing sets. We see each dataset converges to its $Q_{\text{dataset}}$ at a different rate and that there is minimal stochasticity in the theoretical accuracy across replicates.
4.2 Alternatives to PCA
We now compare $Q_{\text{dataset}}$ across different dimensionality reduction schemes. We include a null comparison where we don’t transform the data with a dimensionality reduction scheme (NONE). We compare three linear schemes: PCA, Independent Component Analysis (ICA), and truncated singular vector decomposition / latent semantic analysis (LSA). We also compare three non-linear schemes: non-negative matrix factorization (NMF), radial basis function kernel PCA (KPCA), and uniform manifold approximation and projection (UMAP). All schemes, except UMAP [32], are implemented with Scikit-Learn [33]. Some dimensionality reduction schemes require special processing of the data- for NMF and KPCA, the training and testing sets are min-max normalized before reduction, and for ICA, the number of components is set to the minimum between the number of features and the rank of the training data.
Figure 2: Violin plots of $Q_{\text{dataset}}$ for OpenML datasets across dimensionality reduction schemes. The y-axis is cut-off at $100$ for ease of visualization. Each violin depicts the distribution of $Q_{\text{dataset}}$ for a given scheme by showing its probability density across $Q_{\text{dataset}}$ values.
Figure 2 shows the distributions of $Q_{\text{dataset}}$ across dimensionality reduction schemes. Overall, ICA has the lowest rounded mean $Q_{\text{dataset}}$ at $19$, followed by PCA ($20$), LSA ($21$), KPCA ($21$), NMF ($26$), NONE ($49$), and UMAP ($70$). ICA, PCA, LSA, and KPCA aim to reduce statistical redundancy among features. Specifically, PCA, LSA, and KPCA transform the original feature space into orthogonal components, which are uncorrelated, while ICA produces statistically independent components. If components are statistically dependent, information about the class label $y$ may be redundantly captured across multiple components. To see this, let $X=[x_{1},\dots,x_{D}]$ denote the $D$ components after dimensionality reduction. The true information the components provide about the class is given by the joint mutual information $I(X;y)$ [34]. However, the importance score of a feature $i$ is given by $I(X_{i};y)$, the mutual information between feature $i$ and $y$. Therefore, if the components are statistically dependent, then the sum $\sum_{i=1}^{D}{I(X_{i};y)}$ over-counts overlapping information. For example, if $X_{i}$ and $X_{j}$ are statistical dependent ($I(X_{i};X_{j})>0$), then some portion of the information they provide about $y$ could be redundant. In general, the over-count among all components can be written as: $\sum_{i=1}^{D}{I(X_{i};y)}-I(X;y)$, the overestimation that arises from feature dependence. Since the allocation of bits is based on $\sum_{i}I(X_{i};y)$, bit-bit encoding with statistically dependent components distributes too many bits to redundant features and too few to unique, complementary features. In contrast, if the components are independent, then $\sum_{i=1}^{D}{I(X_{i};y)}=I(X;y)$, so there is no redundancy. Thus, dimensionality reduction schemes that produce independent or de-correlated components can produce smaller $Q_{\text{dataset}}$ if the allocation of bits reflects the true joint information.
4.3 Results on Resource Estimation
Table 1: $Q_{\text{dataset}}$ statistics when the threshold is $0.99$ theoretical accuracy vs $1.0$ for the 55 most-run OpenML datasets that don’t have symbolic features, rounded to the nearest qubit. Each dataset has $10$ replicates with different train/test splits. $Q_{\text{dataset}}$ is calculated with PCA dimensionality reduction.
Dataset Mean $Q_{\text{dataset}}$ $(0.99)$ Mean $Q_{\text{dataset}}(1.0)$
steel-plates-fault 15 20
qsar-biodeg 24 110
wdbc 14 18
phoneme 21 28
diabetes 18 21
The universal approximation property of bit-bit encoding allows us to state that a dataset can be completely modeled with $Q_{\text{dataset}}$ qubits. Hence, we can then say that if $Q_{\text{dataset}}$ is within classical simulation capabilities, $\lessapprox 50$ qubits [21], the corresponding dataset is unlikely to benefit from quantum advantage. Datasets that require $\gtrapprox 50$ qubits are likely promising targets for quantum machine learning within this framework. In Section 4.2, we saw that classical machine learning datasets encoded to minimize statistical dependency between features had $Q_{\text{dataset}}$ of around $20$, on average. These classical machine learning datasets had less than $1,000$ features, less than $10,000$ samples, and less than $10$ classes (Supplementary Figure 1). This result suggests that classic, medium-sized machine learning datasets are unlikely to benefit from quantum advantage. Furthermore, in Section 4.1 we saw a dramatic difference in $Q_{\text{dataset}}$ depending on if the theoretical accuracy threshold at which to record $Q_{\text{train}}$ and $Q_{\text{test}}$ was $0.99$ or $1.0$. In Table 1 we further explore this difference in the top-five most-run OpenML datasets that fulfilled the criteria outlined in Section 4.1. We see, again, a dramatic difference in $Q_{\text{dataset}}$ when the threshold is $0.99$ vs $1.0$, suggesting a trade-off between complete modeling of the training set and generalization on the testing set.
4.4 Schemes for Very Large Datasets
Figure 3: The table shows how the rounded mean $Q_{\text{dataset}}$ of subsamples of the Tahoe dataset scales with the number of features and number of samples.
The default bit-bit encoding algorithm requires the entire dataset to be read into memory, which becomes impractical for large datasets. Furthermore, dimensionality reduction algorithms may not scale well as the dataset size increases, leading to intractable time and memory overhead. For example, PCA dimensionality reduction scales polynomially with dataset size. Learning problems on large datasets are more likely to benefit from quantum advantage, since large datasets should have more bits of information that can only be compressed without collision and overlap with a significant number of qubits. Therefore, we extend the bit-bit encoding algorithm to operate in a batched streaming mode, enabling the processing of large-scale datasets.
Batched bit-bit encoding begins by streaming the training data in batches of a specified size and incrementally fitting the dimensionality reduction algorithm. After all training batches have been processed and the dimensionality reduction model has been fit, the training data are streamed a second time. In this pass, each batch is encoded using the fitted dimensionality reduction scheme, after which batch-level mutual information importance scores are computed and the minimum and maximum values for each feature are recorded. The overall feature importance scores are obtained by averaging the importance scores across all batches. In a final streaming pass, the training data are re-encoded, min–max scaled, and written to disk along with the fitted dimensionality reduction model, feature importance scores, and per-feature minimum and maximum values. The testing data are then streamed and transformed using the saved models and statistics. To compute $Q_{\text{dataset}}$, both training and testing batches are discretized according to the specified number of qubits and the saved feature importance scores. As the training and testing batches are streamed, unique bit strings, the counts of class labels associate with each bit string, and a count of the total number of samples are stored in memory, allowing collision incidences to be computed in roughly the same manner as in the non-batched setting. The main difference is that the test-train collisions are calculated using the most common class label of the test bit string, rather than the true label of the test sample.
We study how $Q_{\text{dataset}}$ scales with both the number of features and the number of samples in a single large dataset. To this end, we subsample from the Tahoe dataset [22], which comprises approximately $100$ million transcriptomic profiles collected across $50$ cancer cell lines and $379$ drugs at varying dosages (corresponding to $1,100$ small-molecule perturbations). The Tahoe-100M dataset [22] is a unique resource created and shared through a collaboration between Vevo Therapeutics and Parse Biosciences’ GigaLab. It was intentionally generated to assist in the development of AI and machine learning tools in the domain of biology. Historically, most developers had to make do with data generated for other purposes which commonly introduces issues such as batch effects and lack of standardized metrics and annotations that accompany large biological datasets created through aggregation from different sources. Containing millions of samples across many cell lines, with each sample having measurements across thousands of interdependent features in response to different compounds delivered at varying dosages, makes this dataset ideal for exploring the applicability and scaling of quantum machine learning algorithms to understanding chaotic, non-linear, high-dimensional biological systems.
In practical terms, this dataset consists of $100$ million samples, each represented by $62,710$ features (gene expression measurements). For our analysis, we restrict the dataset to a single cell line, NCI-H295 (RRID:CVCL_0456), which is the most frequently seen in the Tahoe dataset at $6\%$ of the samples. We formulate the corresponding prediction task as a multiclass classification problem: given the gene expression profile of a sample, identify which drug was applied to the NCI-H295 cell line. After filtering, the resulting Tahoe subset contains $6,040,372$ samples, $52,783$ features, and $65$ classes (the drugs applied to the NCI-H295 cell line, specifically). We then further subsample from the established Tahoe subset so we can investigate how $Q_{\text{dataset}}$ scales with the number of features and number of samples. We subsample the number of samples increasing logarithmically ($100$, $1000$, $10000$) and subsample the number of features logarithmically such that the number of features is less than the number of samples. We replicate each subsample $10$ times with a random train-test split and a biased feature selection. The feature selection is biased such that genes are chosen as features with a probability proportional to their frequency in the Tahoe subset. This is because gene expressions have been found to exhibit a power-law distribution [35], so the biased feature sampling prevent excessive sparsity in the subsample. Samples are subsampled progressively- so, for example, the subsamples with $1000$ samples will contain all the samples from the $100$ sample subsample. We obtain $Q_{\text{dataset}}$ by calculating the training and testing collision incidence when continuously increasing the number of qubits allocated to the subsample in increments of $10$ until the training and testing collision are both 0. We use batched bit-bit encoding to calculate $Q_{\text{dataset}}$, where the batch size is set to the number of samples divided by $10$.
Figure 3 shows the mean $Q_{\text{dataset}}$ over $10$ replicates as the number of features and number of samples increase. A lower number of features with the same number of samples leads to a higher $Q_{\text{dataset}}$ for the Tahoe subsamples. This trend may arise because, with many samples but relatively few features, the samples are densely packed into a smaller feature space, which increases the likelihood of collisions between samples of different classes. To resolve these collisions, the encoding requires greater bit precision per feature, leading to a higher $Q_{\text{dataset}}$. By contrast, when more features are available, samples are more spread out in the feature space, reducing collisions and therefore lowering $Q_{\text{dataset}}$. For all number of samples and number of features we tested, the mean $Q_{\text{dataset}}>50$, suggesting the learning problem on this dataset would benefit from quantum advantage. This finding reflects a growing pattern in which quantum machine learning is emerging as a potentially powerful tool for addressing biological challenges [36, 37].
5 Training
In quantum machine learning, models that are expressive are often associated with barren plateaus and thus thought to be untrainable. While this paper focuses on creating a framework for calculating the number of qubits required for a model, we also show a brief demonstration of training with bit-bit encoding on a dataset consisting of first 4 digits of MNIST. We use the model architecture, sub-net initialization, and exact coordinate update optimization techniques from [19].
We create a restricted training set by removing the collisions at each model size. The training batches consist of only the unique samples that occur after collision removal. This is possible as the number of samples grows slower than exponentially with the number of bits as shown in Fig. 4 (a). Interestingly, even when the dataset is fully encoded, the number of unique samples in the encoded dataset is less than the total number of samples in the dataset.
Figure 4: (left) The number of unique samples as a function of the number of bits used to encode the input. The inset shows the same curve with a log scale on the y-axis and compares it with an exponential. (right) Evolution of test accuracy for the first 4 digits of MNIST. $N_{q}$ is the number of qubits. 2 qubits are reserved for reading out the class label, and $N_{q}-2$ qubits are available to load the data sample at the input.
Figure 4 (right) shows the evolution of test accuracy as a function of the number of training sweeps for models with different numbers of qubits $N_{q}$. In the exact coordinate update optimizer, one parameter is updated at a time, and a sweep denotes one full pass through all parameters. The dashed horizontal lines indicate the theoretical maximum accuracy achievable at that number of qubits. As can be seen, all the models converge to this maximum accuracy predicted by bit-bit encoding. Larger models require more sweeps to converge. They also exhibit transient drops in accuracy due to the use of higher-order loss functions which are more unstable during training.
6 Discussion and Future Directions
In this work, we have examined the universal approximation and resource estimation characteristics of bit-bit encoding for loading classical data into quantum computers. Naively, one may assume that encoding in the basis is useful only when the input data is naturally in the form of binary strings, for example, binary images, combinatorial optimization problems, or molecular orbital occupation vectors. Contrary to this intuition, we have demonstrated that datasets with real-valued entries can also be efficiently represented as bit strings.
We proved that bit-bit encoding inherently ensures universal approximation. Moreover, for a fixed number of qubits, the achievable accuracy can be efficiently determined from the dataset itself, without requiring model training. Taken together, these results establish the first systematic framework for resource estimation in quantum machine learning.
Other encoding schemes have been explored with varying degrees of success, but their expressive power is far less predictable, particularly when classical preprocessing and compression are involved. As a result, they lack a systematic framework for resource estimation and provide no reliable way to judge whether the same outcomes could be achieved with fewer qubits. This makes it nearly impossible to meaningfully assess quantum advantage. While such methods may appear to work on small datasets, without a principled understanding of their expressivity, they cannot be expected to scale. Many previous quantum machine learning studies have also leaned on hybrid models, with part of the model parameterized classically and part quantum, which further blurs the line of advantage. In sharp contrast, our model relies exclusively on quantum parameters, making the presence and scope of quantum advantage in learning far clearer.
An additional advantage of bit-bit encoding is that it enables an intrinsic evaluation of quantum model performance, independent of comparisons to classical baselines, since the maximum achievable accuracy is known a priori. Conventionally, the performance of quantum models is typically judged relative to classical models; however, classical accuracy can vary widely and may even depend on the expertise of the practitioner training the model. By contrast, bit-bit encoding offers an objective criterion for assessing how well a quantum model has learned.
In the numerical sections of the paper, we studied a variety of encoding schemes, and found that the number of qubits they predict are often very close. Therefore, while we cannot theoretically predict that we have an optimal encoding, it is likely to be close to optimal in an information theoretical sense. Following this, we can be more confident in stating that if a dataset can only fit in a number of qubits that surpasses classical simulation capability of around 50 qubits, it will benefit from quantum advantage within this framework. In particular, it is notable that medium-sized OpenML datasets can be fit within 20 qubits using our method and therefore will not go beyond classical simulation capabilities. This is turn indicates that several datasets that have been studied in the quantum literature so far are not candidates for quantum advantage.
An unexpected finding of our work is that a dataset can sometimes be encoded into fewer number of qubits if it has more dimensions. This is in contrast to our initial expectation that a larger dataset would necessarily require more quantum resources to model. It indicates that the use of the mutual information in the encoding algorithm allows for the efficient extraction of information relevant to the learning task from the original dataset. This phenomenon, to our knowledge, cannot occur with alternative encoding schemes.
For future work, we plan to extend bit-bit encoding to convert more data types into quantum basis states, including sequential, multi-label, regression, and spatial. We also plan to focus on investigating the potential for quantum advantage for datasets from specific fields, such as biology. We will aid this investigation by developing information-theoretic metrics that may support prediction of $Q_{\text{dataset}}$ without iterating through the whole range of qubits, and further validate $Q_{\text{dataset}}$ with training experiments on a variety of datasets.
While this work has primarily focused on determining the number of qubits required for encoding, many open directions remain in developing a deeper understanding of bit-bit encoding. One important question concerns the optimal number of qubits for generalization: although increasing qubits improves precision, it may also introduce a trade-off between perfectly fitting the training data and effectively generalizing to unseen examples. Another open challenge is to move beyond logical qubits and identify what ansatz structures yield efficient models when implemented on physical qubits. Finally, questions remain about the gate complexity and learning dynamics of models trained on bit-bit encoded data. These issues will be the focus of future investigations.
7 Software Framework
The techniques in the paper are implemented in Red Cedar, a software framework being developed at Coherent Computing Inc. It can be made available upon request.
Supplemental Figures
Figure 5: The number of features (x, log-scale), number of samples (y, log-scale), and number of classes (hue) of each of the $29$ OpenML datasets used in experiments for Section 4.1 and Section 4.2.
Figure 6: The theoretical training and testing accuracy across allocated qubits for the experiment with PCA dimensionality reduction. Each line is the mean for each dataset, with $95\%$ confidence interval error bands.